OpenStudy (anonymous):

identify the slope and the y-intercept of the equation y = -1/3x + 3 Describe how you would graph this line using the slope-intercept method.

OpenStudy (harsimran_hs4):

y = mx +c is the equation of straight line where m = slope and c = intercept on the y axis (basic known result) can you figure out now?

OpenStudy (anonymous):

so -1/3 is the slope? and 3 is the y-intercept?

OpenStudy (e.cociuba):

slope= -1/3x y-intercept=3

OpenStudy (anonymous):

but how would you graph it using the slope-intercept method?

OpenStudy (e.cociuba):

you go up to (0,3), and then go up 1 and to the left 3

OpenStudy (e.cociuba):

do u get it??

OpenStudy (anonymous):

so to make it into a coordinate you just make 0 your x? if so then i get it :)

OpenStudy (e.cociuba):

yes im pretty sure:)

OpenStudy (anonymous):

thank you soso much!

OpenStudy (e.cociuba):

np:)